# Classification of connection graphs of global attractors for $S^{1}$-equivariant parabolic equations

### v2

@@ 2 Homotopy between $f_{0}(u,u_{x})$ and $f_{1}(u)$
For references see Theorem 4.2.5 and Proposition 4.2.6 of [32], Theorem 4.2 of [14], and [31].
@@
(2.1) $f(u,u_{x})=\frac{1}{k}u\left(ku^{2}+\sqrt{k^{2}u^{4}+4u_{x}^{2}}\right)\ .$
Each level curve of the first integral $I(u,v)=\frac{k}{2}u\left(ku^{2}+\sqrt{k^{2}u^{4}+4v^{2}}\right)$ of (1.2) is an ellipse $v^{2}+kI_{0}u^{2}=I_{0}$ where $I_{0}:=I(u_{0},0)=ku_{0}^{2}$. Terefore, the period map for this nonlinearity is given by
@@
If the period map $T(u_{0})$ satisfies condition (C) we are able to find a Hamiltonian nonlinearity $g(u)$ with the same period map, [30]. See also [33, 32] for the preliminary results. In this case, the convex combination
@@
realizes the desired homotopy in the Hamiltonian class of nonlinearities.
On the other hand, the failing of condition (C) prevents the homotopy realization in the Hamiltonian class since the period map is not realizable in this class. Therefore, the homotopy has to be constructed directly in $\mathcal{R}$ using the phase portrait of (1.2). We define the homotopy as a continuous transformation of the cyclicity set $\mathcal{C}$ of $f(u,u_{x})$, which preserves the hyperbolicity of all homogeneous equilibria and frozen rotating waves, and shrinks the period map along the $u$-axis until it satisfies condition (C).
Proof of Theorem 1: First, let $f_{1}(u,u_{x})$ denote a reversible nonlinearity such that its time map $T=T(u_{0})$ is realizable by a Hamiltonian nonlinearity $g=g(u)$. Also, let $G$ denote the potential function of (1.2), i.e. $G^{\prime}=g$. Under the hyperbolicity assumption (H) for (1.1) with $f=f_{1}(u,u_{x})$, all the zeros of $g$ are nondegenerate (i.e. $g(u_{0})=0$ implies $g^{\prime}(u_{0})\neq 0$). Moreover, the period map $T=T(u_{0})$ satisfies the nondegeneracy condition: $T(u_{0})=\frac{2\pi}{k}$ implies $T^{\prime}(u_{0})\neq 0$ for all positive integers $k\in\mathbb{N}$. See [30] for details.
@@
Initially, we consider that the lap signature class of $f(u,u_{x})$ is of simple type. This implies that $l=0$ and we deal with a cyclicity set composed only of isolated punctured disks around the centers in the phase portrait.
Figure 1: Cyclicity set $\mathcal{C}_{0}$ in the case of $n=3,l=0$ and with $q=2$ $2\pi$-periodic orbits.
The proof essentially consists on the construction of a homotopy in $\mathcal{R}$ between $f_{0}(u,u_{x})$ and a nonlinearity $f_{1}(u,u_{x})$ satisfying condition (C) of a Hamiltonian nonlinearity $g=g_{1}(u)$. This homotopy preserves: (a) hyperbolicity of all critical elements of the flow generated by (1.1) with $f=f_{0}(u,u_{x})$; and hence, (b) the lap signature of the period map $T=T_{f}$.
@@
Case (I): $n=3$ and $l=0$. For the reversible nonlinearity $f=f_{0}(u,u_{x})$ let $e_{1}<e_{2}=0<e_{3}$ denote the three zeros of $f_{0}(\cdot,0)$ corresponding to two saddle points, ${\bf e}_{1},{\bf e}_{3}$, and a center at the origin ${\bf e}_{2}=(0,0)$. The boundary $\partial\mathcal{C}_{0}$ of the cyclicity set is given by a saddle point and an orbit homoclinic to this saddle. Without loss of generality we assume that the saddle point is ${\bf e}_{1}$. Therefore, the cyclicity set $\mathcal{C}_{0}$ is positioned to the right of ${\bf e}_{1}$, see Fig. 1.
@@
To define the scale function $\omega|(0,a^{+})$, let $c_{1},c_{2}\in(\alpha_{1},a^{+})$ denote two constants satisfying $\alpha_{1}<c_{1}<c_{2}<a^{+}$, where again $\alpha_{1}$ denotes the maximal value of the outermost $2\pi$-periodic solution $v_{1}(\cdot,a_{1})$. Recall that $\lim_{a\rightarrow a^{+}}T(a)=+\infty$, hence $T(a)>2\pi$ for $a\in(\alpha_{1},a^{+})$, see Fig. 2. Moreover, $T^{\prime}(a)>0$ in a neighborhood of $a^{+}$ and we always choose $c_{2}$ in this neighborhood. Then, for a parameter $\delta\in(0,1)$, we define
@@
Therefore, with this scale function $\omega$, the diffeomorphism $\Phi$ shrinks all the $2\pi$-periodic orbits to a neighborhood of the origin as $\delta$ decreases, see Fig. 3. Moreover, the period map $T_{1}(a)$ satisfies
@@
(2.9) $\frac{d}{da}\left(aT_{1}(a)\right)=T(\omega(a)a)+\omega(a)aT^{\prime}(\omega(a)a)\ ,$
which implies that, for $\delta$ sufficiently small, $T_{1}(a)$ satisfies condition (C). Notice that by our choice of $c_{2}$ in a neighborhood of $a^{+}$ the period map satisfies $T^{\prime}_{1}(a)>0$. We conclude that $f_{1}(u,u_{x})$ is realizable in the Hamiltonian class of nonlinearities completing the proof of case (I).
Case (II): $n\geq 5$ and $l=0$. Now we deal with a cyclicity set
(2.10) $\mathcal{C}=\bigcup_{1\leq k\leq(n+1)/2}\mathcal{C}_{k}$
composed of several punctured disks surrounding the $(n+2)/2$ centers of (1.2), see Fig. 4. Since each $\mathcal{C}_{k}$ is isolated (by condition (M) and the simple type $l=0$), the proof in this case follows by repeating the previous procedure in each region $\mathcal{C}_{k},k=1,\dots,(n+1)/2$. Therefore, we obtain $(n+1)/2$ disjoint graphs of Hamiltonian realizations, one for each region $\mathcal{C}_{k}$. Then, a Hamiltonian $g_{1}(u)$ which realizes $f_{1}(u,u_{x})$ is obtained by extending the domain to $\mathbb{R}$ joining these $g_{1}$ graphs. For this extension, we fill the $(n-3)/2$ closed interval gaps with $C^{2}$-smooth functions without introducing further zeros of $g_{1}$. Similarly, in the two remaining unbounded intervals, our extension choice is such that no further zeros are introduced. Finally, we choose a globally bounded extension $g_{1}:\mathbb{R}\rightarrow\mathbb{R}$.
Figure 4: Cyclicity set $\mathcal{C}=\mathcal{C}_{1}\cup\mathcal{C}_{2}$ for $f=f_{0}(u,u_{x})$ with $n=5$ and $l=0$.
Clearly, $f_{0}(u,u_{x})$ and $g_{1}(u)$ belong to the same lap signature class. The desired homotopy, then, has the form (see [31]),
@@
followed by the homotopy (2.3) between $f_{1}(u,u_{x})$ and $g_{1}(u)$. Finally, as a result of [31], Theorem 2, (see also [30, 12]), there exists a homotopy between $f(u)$ and $g(u)$ since they belong to the same lap signature class. This completes the proof of case (II).
Remark: Consider again the Hamiltonian realization of $f=f_{0}(u,u_{x})$ by $g_{1}(u)$ in case (I). The cyclicity set $\mathcal{C}_{0}$ is right oriented if its boundary $\partial\mathcal{C}_{0}$ contains the saddle point ${\bf e}_{1}$, i.e. $a^{-}=e_{1}$. Similarly, $\mathcal{C}_{0}$ is left oriented if $\partial\mathcal{C}_{0}$ contains the saddle point ${\bf e}_{3}$, i.e. $a^{+}=e_{3}$. Then, if $\mathcal{C}_{0}$ is right oriented, the nonlinearity $g_{1}$ satisfies $g_{1}(a^{-})=g_{1}(e_{1})=0$ and is negative for $u\in(a^{-},0)$ and positive for $u\in(0,a^{+})$. Therefore, the extension of $g_{1}$ to the interval gap $(a^{+},e_{3})$ is also positive. On the other hand, if $\mathcal{C}_{0}$ is left oriented, we have $g_{1}(e_{2})=g_{1}(a^{+})=0$ and the extension of $g_{1}$ to the interval gap $(e_{1},a^{-})$ is negative. This holds in the multiple case (II) for all the isolated components $\mathcal{C}_{k}$ of the cyclicity set.
@@
Figure 5: Cyclicity set $\mathcal{C}=\mathcal{C}_{1}\cup\mathcal{C}_{2}$ for $f=f_{0}(u,u_{x})$ with $n=5$, $l=0$ and $m=3$. The previous Fig. 4 illustrates the case $m=n=5$.
The phase portrait of (1.2) for $f=g_{1}(u)$ depends essentially on the singular values of the potential function
@@
at the saddle points, i.e. the local maxima of $G$, and not on the Morse type of $G$. Indeed, the Morse type of $G$ is easily modified by changing the values of the extended $g_{1}$ in its $(n-1)/2$ closed interval gaps.
Since all $\mathcal{C}_{k}$ are right oriented at the left of ${\bf e}_{m}$ and left oriented at the right, all the extensions of $g_{1}$ are positive to the left of $e_{m}$ and negative to the right. Then, by (2.13), ${\bf e}_{m}$ is the saddle point with the maximum value $G(e_{m})=\max\{G(e_{2k-1}):1\leq k\leq(n+1)/2\}$ and the sequence of singular values $G(e_{2k-1})$ satisfies the inequalities
@@
This ensures that the phase portraits of (1.2) for $f=f_{0}(u,u_{x})$ and $f=g_{1}(u)$ are qualitatively the same.
@@
In view of the proof of previous cases, we only need to consider nonlinearities $f_{0}(u,u_{x})$ of non-simple type, $l\geq 1$. This implies $n\geq 5$. As before we construct a homotopy in $\mathcal{R}$ between $f_{0}(u,u_{x})$ and a nonlinearity $f_{1}(u,u_{x})$ satisfying condition (C).
Case (III): $n=5$ and $l=1$. In this case, the phase portrait of (1.2) for $f=f_{0}(u,u_{x})$ has three cyclicity regions. In addition to two punctured disks $\mathcal{C}_{2},\mathcal{C}_{3}$ surrounding the centers ${\bf e}_{2},{\bf e}_{4}$, there is an annular region $\mathcal{C}_{1}$ surrounding both punctured disks. This implies that the saddle point ${\bf e}_{3}$ has two homoclinic orbits composing the boundary $\partial(\mathcal{C}_{2}\cup\mathcal{C}_{3})$. The cyclicity regions $\mathcal{C}_{2}$ and $\mathcal{C}_{3}$ have opposite left/right orientations and the homoclinic orbits form an $\infty$ shaped curve. See Fig. 6.
Figure 6: Cyclicity set $\mathcal{C}=\mathcal{C}_{1}\cup\mathcal{C}_{2}\cup\mathcal{C}_{3}$ for $f=f_{0}(u,u_{x})$ with $n=5$, $l=1$ and $q=4$. Two of these $2\pi$-periodic orbits are in the annular region $\mathcal{C}_{1}$.
@@
Then, let again $c_{1},c_{2}$, denote the two constants satisfying $\alpha_{1}<c_{1}<c_{2}<a^{+}$ ($c_{2}$ in the appropriate neighborhood of $a^{+}$) and, for $\delta\in(0,1)$, define the scale function $\omega$ by (2.7).
Figure 7: Upper left: Cyclicity set $\mathcal{C}$ for $f=f_{0}(u,u_{x})$. Upper right: Cyclicity set $\Phi\mathcal{C}$ for $f=f_{1}(u,u_{x})$. Bottom: Graph of the scale function $\omega$. Here $b^{-}$ and $b^{+}$ denote respectively the minimum and maximum $u$-values of the orbits homoclinic to the origin. The white balls correspond to these homoclinic orbits, and $b^{\pm}\not\in\mathcal{D}$ since the homoclinic orbits are not in $\mathcal{C}$.
Once more, using this scale function $\omega$, we define the diffeomorphism $\Phi:\mathcal{C}\rightarrow\mathcal{C}$ which we extend to $\mathop{\mathrm{cl}}\nolimits\mathcal{C}$ by continuity, and to the phase plane $\mathbb{R}^{2}$ by the identity, $\Phi|(\mathbb{R}^{2}\setminus\mathop{\mathrm{cl}}\nolimits\mathcal{C})=\mathop{\mathrm{id}}\nolimits$.
As expected, $\Phi$ shrinks all equilibria and $2\pi$-periodic orbits in $\mathcal{C}$ to a neighborhood of the origin $(e_{3},0)=(0,0)$ as $\delta$ decreases, preserving hyperbolicity, see Fig. 7. In addition, under $\Phi$ the period map $T_{1}$ of (1.2) for $f=f_{0}(u,u_{x})$ satisfies (2.8). Then, from (2.9) we again conclude that $T_{1}(a)$ satisfies condition (C). Hence $f_{1}(u,u_{x})$ is realizable by a Hamiltonian nonlinearity $g_{1}(u)$ and the desired homotopy from $f_{1}(u,u_{x})$ to $g_{1}(u)$ has the form (2.11), (2.12), which is followed by the homotopy (2.3). This completes the proof in case (III).
Case (IV): $n\geq 7$ and $l\geq 1$. We proceed sequentially, following the total order imposed by the regular parenthesis structure of the cyclicity regions. For example, the regular parenthesis structure of the cyclicity region shown in Fig. 6 is
@@
The use of the regular parenthesis structure of the cyclicity regions instead of the $2\pi$-periodic orbits is necessary to overcome the case of $l\geq 1$ without $2\pi$-periodic orbits in some annular region $\mathcal{C}_{k}$, i.e. $q_{k}=0$.
@@
We let $\mathcal{C}$ denote the union of $\widetilde{\mathcal{C}}$ with all the cyclicity regions that it encloses. Then, we apply the same argument used in case (III) shrinking all equilibria and $2\pi$-periodic orbits contained in $\mathcal{C}$ to a neighborhood of the unique saddle point in the inner boundary of $\widetilde{\mathcal{C}}$. For example, in Fig. 6 the saddle point is ${\bf e}_{3}$. Therefore, using the same notation as in case (III), we obtain a diffeomorphism $\Phi:\mathcal{C}\rightarrow\mathcal{C}$ which preserves hyperbolicity and leads to a period map $T_{1}|(a^{-},a^{+})$ satisfying condition (C).
@@
Hence, we again extend $\Phi$ to $\mathop{\mathrm{cl}}\nolimits\mathcal{C}$ by continuity and to the complete phase space $(u,u_{x})\in\mathbb{R}^{2}$ by the identity $\Phi|(\mathbb{R}^{2}\setminus\mathop{\mathrm{cl}}\nolimits\mathcal{C})=\mathop{\mathrm{id}}\nolimits$. Notice that each $\mathcal{C}_{k}$ is isolated and, due to this isolation and the integrability of (1.1) for $f=f_{1}(u,u_{x})$, each $\mathcal{C}_{k}$ accepts the left/right orientation described in the previous Remark after case (II).
We obtain a Hamiltonian realization of (1.2) for $f=f_{1}(u,u_{x})$ by a nonlinearity $g_{1}(u)$ defined on the isolated intervals determined by the regions $\mathcal{C}_{k}$. Then, after the $C^{2}$ smooth and globally bounded extension of $g_{1}(u)$ to $\mathbb{R}$, filling the unbounded intervals and the interval gaps, we define the potential $G$ by (2.13). Therefore, due to the partial order (2.14) restricted to the saddle points on the boundaries of $\mathcal{C}_{k}$, the phase portraits of (1.2) for $f=f_{0}(u,u_{x})$ and $f=g_{1}(u)$ are qualitatively the same.

### v5

@@ 3 Homotopy construction
For references see Theorem 4.2.5 and Proposition 4.2.6 of [36], Theorem 5 of [37], Theorem 4.2 of [17], and Section 4 of [34].
@@
(3.1) $f(u,u_{x})=\frac{k}{2}u\left(ku^{2}+\sqrt{k^{2}u^{4}+4u_{x}^{2}}\right)\ .$
Each level curve of the first integral $I(u,v)=\frac{1}{2}\left(ku^{2}+\sqrt{k^{2}u^{4}+4v^{2}}\right)$ of (1.2) is an ellipse $v^{2}+kI_{0}u^{2}=I_{0}^{2}$ where $I_{0}:=I(u_{0},0)=ku_{0}^{2}$. Therefore, the period map for this nonlinearity is given by
@@
If the period map $T(u_{0})$ satisfies condition (C) we are able to find a Hamiltonian pendulum nonlinearity $g=g(u)$ with the same period map, [34]. See also [37, 36] for the preliminary results. In this case, the convex combination
@@
provides the desired pendulum realization homotopy in the Hamiltonian class of nonlinearities.
On the other hand, the failing of condition (C) prevents the homotopy in the Hamiltonian class since the period map is not realizable in this class. Therefore, the homotopy has to be constructed directly in $\mathcal{R}$ using the phase portrait of (1.2). We define the Hamiltonian realization homotopy as a continuous transformation of the cyclicity set $\mathcal{C}$ of $f(u,u_{x})$, which preserves the hyperbolicity of all homogeneous equilibria and frozen rotating waves, and shrinks the period map along the $u$-axis until it satisfies condition (C).
Let $K=(n-1)/2+r$ denote the number of connected regions of the cyclicity set $\mathcal{C}$. Here $n$ denotes the odd number of equilibria and $r$ the number of annular regions. Annular regions are the multiply connected regions bounded by two saddles and their homoclinic ODE separatrices (see for example the region $\mathcal{C}_{1}$ in Fig.6). Notice that each equilibrium of (1.1) corresponds to a center or a saddle point of the phase portrait of (1.2). The centers and saddles alternate for increasing values of $u_{0}$, starting and ending with saddles. Then $(n-1)/2$ is the number of punctured disks encircling the centers. Moreover, $r$ is the number of annular regions which surround more than one single center. This implies that $20\leq r\leq(n-3)/$ 2. See [15].
Finally, a reversible nonlinearity $f=f(u,u_{x})$ is of simple type if $r=0$, i.e. if there are no annular regions. This is a slightly more restrictive definition then the one used in [35]. If $r=0$ then each $2\pi$-periodic orbit of (1.2) encircles exactly one center in the phase plane $(v,v_{x})$.
Then, let $g=g(u)$ denote a Hamiltonian nonlinearity with period map $T=T(u_{0})$, and let $G$ denote the potential function of (1.2), i.e. $G^{\prime}=g$. Under the hyperbolicity assumption (H) for (1.1) with $f=f_{1}(u,u_{x})$, all the zeros of $g$ are nondegenerate (i.e. $g(u_{0})=0$ implies $g^{\prime}(u_{0})\neq 0$). Moreover, the period map $T=T(u_{0})$ satisfies the nondegeneracy condition: $T(u_{0})=\frac{2\pi}{k}$ implies $T^{\prime}(u_{0})\neq 0$ for all positive integers $k\in\mathbb{N}$. See [34] for details.
@@
Initially, we consider that the lap signature class of $f(u,u_{x})$ is of simple type. This implies that $r=0$ and we deal with a cyclicity set composed only of isolated punctured disks around the centers in the phase portrait.
Figure 1: Cyclicity set $\mathcal{C}_{0}$ in the case of $n=3,r=0$ and with $q=2$ $2\pi$-periodic orbits.
Proof of Theorem 2: The proof essentially consists on the construction of a Hamiltonian realization homotopy in $\mathcal{R}$ between $f_{0}(u,u_{x})$ and a nonlinearity $f_{1}(u,u_{x})$ satisfying condition (C) of a Hamiltonian nonlinearity $g=g(u)$. This homotopy preserves: (a) hyperbolicity of all critical elements of the flow generated by (1.1) with $f=f_{0}(u,u_{x})$; and hence, (b) the lap signature of the period map $T=T_{f}$.
@@
Case (I): $n=3$ and $r=0$. For the reversible nonlinearity $f=f_{0}(u,u_{x})$ let $e_{1}<e_{2}=0<e_{3}$ denote the three zeros of $f_{0}(\cdot,0)$ corresponding to two saddle points, ${\bf e}_{1},{\bf e}_{3}$, and a center at the origin ${\bf e}_{2}=(0,0)$. The boundary $\partial\mathcal{C}_{0}$ of the cyclicity set is given by a saddle point and an orbit homoclinic to this saddle. Without loss of generality we assume that the saddle point is ${\bf e}_{1}$. Therefore, the cyclicity set $\mathcal{C}_{0}$ is positioned to the right of ${\bf e}_{1}$, see Fig. 1.
@@
To define the scale function $\omega|_{(0,a^{+})}$, let $c_{1},c_{2}\in(\alpha_{1},a^{+})$ denote two constants satisfying $\alpha_{1}<c_{1}<c_{2}<a^{+}$, where again $\alpha_{1}$ denotes the maximal value of the outermost $2\pi$-periodic solution $v_{1}(\cdot,a_{1})$. Recall that $\lim_{a\rightarrow a^{+}}T(a)=+\infty$, hence $T(a)>2\pi$ for $a\in(\alpha_{1},a^{+})$, see Fig. 2. Changing the phase portrait of (1.2) in a small neighborhood of the homoclinic orbit to ${\bf e}_{1}$ does not affect the global attractor $\mathcal{A}_{f}$. Therefore, we can assume that the period map satisfies $\lim_{a\rightarrow a^{+}}T(a)=+\infty$ monotonically. Hence, we have $T^{\prime}(a)>0$ in a neighborhood of $a^{+}$ and we always choose $c_{1}$ in this neighborhood. Then, for a parameter $\delta\in(0,1)$, we define
@@
Therefore, with this scale function $\omega$, the diffeomorphism $\Phi$ shrinks all the $2\pi$-periodic orbits to a neighborhood of the origin as $\delta$ decreases, see Fig. 3. Since the scaling of the period map preserves the nondegeneracy of the points corresponding to the maximal values of the $2\pi$-periodic solutions, their hyperbolicity is preserved by the scaling. Moreover, the period map $T_{1}(a)$ satisfies
@@
(3.9) $\frac{d}{da}\left(aT_{1}(a)\right)=T(\omega(a)a)+(\omega(a)+a\omega^{\prime}(a))aT^{\prime}(\omega(a)a)\ .$
Notice that $\omega^{\prime}(a)=0$ for all $a>0$ except $a\in(c_{1},c_{2})$ where $\omega^{\prime}(a)>0$ and, by our choice of $c_{1}$ in a neighborhood of $a^{+}$, the period map satisfies $T^{\prime}(\omega(a)a)>0$ there. This implies that, for $\delta$ sufficiently small, $T_{1}(a)$ satisfies condition (C). We conclude that $f_{1}(u,u_{x})$ is realizable in the Hamiltonian class of nonlinearities completing the proof of case (I).
Case (II): $n\geq 5$ and $r=0$. Now we deal with a cyclicity set
(3.10) $\mathcal{C}=\bigcup_{1\leq k\leq(n-1)/2}\mathcal{C}_{k}$
composed of several punctured disks surrounding the $(n-1)/2$ centers of (1.2), see Fig. 4. Since each $\mathcal{C}_{k}$ is isolated (by condition (M) and the simple type $r=0$), the proof in this case follows by repeating the previous procedure in each region $\mathcal{C}_{k},k=1,\dots,(n-1)/2$. Therefore, we obtain $(n-1)/2$ disjoint graphs of Hamiltonian realizations, one for each region $\mathcal{C}_{k}$. Then, a Hamiltonian $g(u)$ which realizes $f_{1}(u,u_{x})$ is obtained by extending the domain to $\mathbb{R}$ joining these $g$ graphs. For this extension, we fill the $(n-3)/2$ closed interval gaps with $C^{2}$-smooth functions without introducing further zeros of $g$. Similarly, in the two remaining unbounded intervals, our extension choice is such that no further zeros are introduced. Finally, we choose a globally bounded extension $g:\mathbb{R}\rightarrow\mathbb{R}$.
Figure 4: Cyclicity set $\mathcal{C}=\mathcal{C}_{1}\cup\mathcal{C}_{2}$ for $f=f_{0}(u,u_{x})$ with $n=5$ and $r=0$.
Clearly, $f_{0}(u,u_{x})$ and $g(u)$ belong to the same lap signature class. Then, the desired Hamiltonean realization homotopy has the form (see [35]),
@@
followed by the pendulum realization homotopy (3.3) between $f_{1}(u,u_{x})$ and $g(u)$. This completes the proof of case (II).
Remark: Consider again the Hamiltonian realization of $f=f_{0}(u,u_{x})$ by $g(u)$ in case (I). The cyclicity set $\mathcal{C}_{0}$ is right oriented if its boundary $\partial\mathcal{C}_{0}$ contains the saddle point ${\bf e}_{1}$, i.e. $a^{-}=e_{1}$. Similarly, $\mathcal{C}_{0}$ is left oriented if $\partial\mathcal{C}_{0}$ contains the saddle point ${\bf e}_{3}$, i.e. $a^{+}=e_{3}$. Then, if $\mathcal{C}_{0}$ is right oriented, the nonlinearity $g$ satisfies $g(a^{-})=g(e_{1})=0$ and is negative for $u\in(a^{-},0)$ and positive for $u\in(0,a^{+})$. Therefore, the extension of $g$ to the interval gap $(a^{+},e_{3})$ is also positive. On the other hand, if $\mathcal{C}_{0}$ is left oriented, we have $g(e_{2})=g(a^{+})=0$ and the extension of $g$ to the interval gap $(e_{1},a^{-})$ is negative. This holds in the multiple case (II) for all the isolated components $\mathcal{C}_{k}$ of the cyclicity set.
@@
Figure 5: Cyclicity set $\mathcal{C}=\mathcal{C}_{1}\cup\mathcal{C}_{2}$ for $f=f_{0}(u,u_{x})$ with $n=5$, $r=0$ and $m=3$. The previous Fig. 4 illustrates the case $m=n=5$.
The phase portrait of (1.2) for $f=g(u)$ depends essentially on the singular values of the potential function
@@
at the saddle points, i.e. the local maxima of $G$, and not on the Morse type of $G$. Indeed, the Morse type of $G$ is easily modified by changing the values of the extended $g$ in its $(n-1)/2$ closed interval gaps.
Since all $\mathcal{C}_{k}$ are right oriented at the left of ${\bf e}_{m}$ and left oriented at the right, all the extensions of $g$ are positive to the left of $e_{m}$ and negative to the right. Then, by (3.13), ${\bf e}_{m}$ is the saddle point with the maximum value $G(e_{m})=\max\{G(e_{2k-1}):1\leq k\leq(n+1)/2\}$ and the sequence of singular values $G(e_{2k-1})$ satisfies the inequalities
@@
This ensures that the phase portraits of (1.2) for $f=f_{0}(u,u_{x})$ and $f=g(u)$ are qualitatively the same.
@@
In view of the proof of previous cases, we only need to consider nonlinearities $f_{0}(u,u_{x})$ of non-simple type, $r\geq 1$. This implies $n\geq 5$. As before we construct a Hamiltonian realization homotopy in $\mathcal{R}$ between $f_{0}(u,u_{x})$ and a nonlinearity $f_{1}(u,u_{x})$ satisfying condition (C).
Case (III): $n=5$ and $r=1$. In this case, the phase portrait of (1.2) for $f=f_{0}(u,u_{x})$ has three cyclicity regions. In addition to two punctured disks $\mathcal{C}_{2},\mathcal{C}_{3}$ surrounding the centers ${\bf e}_{2},{\bf e}_{4}$, there is an annular region $\mathcal{C}_{1}$ surrounding both punctured disks. This implies that the saddle point ${\bf e}_{3}$ has two homoclinic orbits composing the boundary $\partial(\mathcal{C}_{2}\cup\mathcal{C}_{3})$. The cyclicity regions $\mathcal{C}_{2}$ and $\mathcal{C}_{3}$ have opposite left/right orientations and the homoclinic orbits form an $\infty$ shaped curve. See Fig. 6.
Figure 6: Cyclicity set $\mathcal{C}=\mathcal{C}_{1}\cup\mathcal{C}_{2}\cup\mathcal{C}_{3}$ for $f=f_{0}(u,u_{x})$ with $n=5$, $r=1$ and $q=4$. Two of these $2\pi$-periodic orbits are in the annular region $\mathcal{C}_{1}$.
@@
Then, let again $c_{1},c_{2}$, denote the two constants satisfying $\alpha_{1}<c_{1}<c_{2}<a^{+}$ ($c_{1}$ in the appropriate neighborhood of $a^{+}$) and, for $\delta\in(0,1)$, define the scale function $\omega$ by (3.7).
Figure 7: Upper left: Cyclicity set $\mathcal{C}$ for $f=f_{0}(u,u_{x})$. Upper right: Cyclicity set $\Phi\mathcal{C}$ for $f=f_{1}(u,u_{x})$. Bottom: Graph of the scale function $\omega$. Here $b^{-}$ and $b^{+}$ denote respectively the minimum and maximum $u$-values of the orbits homoclinic to the origin. The white balls correspond to these homoclinic orbits, and $b^{\pm}\not\in\mathcal{D}$ since the homoclinic orbits are not in $\mathcal{C}$. Also note that $0\not\in\mathcal{D}$.
Once more, using this scale function $\omega$, we define the diffeomorphism $\Phi:\mathcal{C}\rightarrow\mathcal{C}$ which we extend to $\mathop{\mathrm{cl}}\nolimits\mathcal{C}$ by continuity, and to the phase plane $\mathbb{R}^{2}$ by the identity, $\Phi|_{(\mathbb{R}^{2}\setminus\mathop{\mathrm{cl}}\nolimits\mathcal{C})}=\mathop{\mathrm{id}}\nolimits$.
As expected, $\Phi$ shrinks all equilibria and $2\pi$-periodic orbits in $\mathcal{C}$ to a neighborhood of the origin $(e_{3},0)=(0,0)$ as $\delta$ decreases, preserving hyperbolicity, see Fig. 7. In addition, under $\Phi$ the period map $T_{1}$ obtained from the period map $T$ of (1.2) for $f=f_{0}(u,u_{x})$ satisfies (3.8). Then, from (3.9) we again conclude that $T_{1}(a)$ satisfies condition (C). Hence $f_{1}(u,u_{x})$ is realizable by a Hamiltonian nonlinearity $g(u)$ and the desired homotopy from $f_{1}(u,u_{x})$ to $g(u)$ has the form (3.11), (3.12), which is followed by the pendulum realization homotopy (3.3). This completes the proof in case (III).
Case (IV): $n\geq 7$ and $r\geq 1$. We proceed sequentially, following the total order imposed by the regular parenthesis structure of the cyclicity regions. For example, the regular parenthesis structure of the cyclicity region shown in Fig. 6 is
@@
The use of the regular parenthesis structure of the cyclicity regions instead of the $2\pi$-periodic orbits is necessary to overcome the case of $r\geq 1$ without $2\pi$-periodic orbits in some annular region $\mathcal{C}_{k}$, i.e. $q_{k}=0$.
@@
We let $\mathcal{C}$ denote the union of $\widetilde{\mathcal{C}}$ with all the cyclicity regions that it encloses. Then, we apply the same argument used in case (III) shrinking all equilibria and $2\pi$-periodic orbits contained in $\mathcal{C}$ to a neighborhood of the unique saddle point in the inner boundary of $\widetilde{\mathcal{C}}$. For example, in Fig. 6 the saddle point is ${\bf e}_{3}$. Therefore, using the same notation as in case (III), we obtain a diffeomorphism $\Phi:\mathcal{C}\rightarrow\mathcal{C}$ which preserves hyperbolicity and leads to a period map $T_{1}|_{(a^{-},a^{+})}$ satisfying condition (C).
@@
Hence, we again extend $\Phi$ to $\mathop{\mathrm{cl}}\nolimits\mathcal{C}$ by continuity and to the complete phase space $(u,u_{x})\in\mathbb{R}^{2}$ by the identity $\Phi|_{(\mathbb{R}^{2}\setminus\mathop{\mathrm{cl}}\nolimits\mathcal{C})}=\mathop{\mathrm{id}}\nolimits$. Notice that each $\mathcal{C}_{k}$ is isolated and, due to this isolation and the integrability of (1.1) for $f=f_{1}(u,u_{x})$, each $\mathcal{C}_{k}$ accepts the left/right orientation described in the previous Remark after case (II).
We obtain a Hamiltonian and pendulum realization of (1.2) for $f=f_{1}(u,u_{x})$ by a nonlinearity $g(u)$ defined on the isolated intervals determined by the regions $\mathcal{C}_{k}$. Then, after the $C^{2}$ smooth and globally bounded extension of $g(u)$ to $\mathbb{R}$, filling the unbounded intervals and the interval gaps, we define the potential $G$ by (3.13). Therefore, due to the partial order (3.14) restricted to the saddle points on the boundaries of $\mathcal{C}_{k}$, the phase portraits of (1.2) for $f=f_{0}(u,u_{x})$ and $f=g(u)$ are qualitatively the same. Then, preservation of hyperbolicity ensures that $f=f_{0}(u,u_{x})$ and $f=g(u)$ belong to the same lap signature class. This completes the proof of case (IV).
Finally, as a result of [35], Theorem 2, (see also [34, 15]), there exists a homotopy between the pendulum realization $g(u)$ of $f_{0}(u,u_{x})$ and a pendulum realization $\overline{g}(u)$ of $f_{1}(u,u_{x})$ since, by assumption, both belong to the same lap signature class. In fact, all global attractors of pendulum realizations with the same lap signature are connection equivalent.
In summary, the global collective homotopy is the composition of:
  1. (i)
    a Hamiltonian realization homotopy (3.11), (3.12), and a pendulum realization homotopy (3.3) from $f_{0}(u,u_{x})$ to $g(u)$;
  2. (ii)
    a homotopy between $g(u)$ and $\overline{g}(u)$;
  3. (iii)
    a reverse pendulum realization homotopy and a reverse Hamiltonian realization homotopy from $\overline{g}(u)$ to $f_{1}(u,u_{x})$.